OpenStudy (anonymous):

confused plz help.. Water freezes at 0degrees Celsius and 32degrees Fahrenheit. Water boils at 100degrees Celsius and 212degrees Fahrenheit. a. Write degrees Celsius as a linear function of degrees Fahrenheit b. What is the slope of your linear equation? What does it mean? c. What is the y-intercept and what does it mean?

OpenStudy (phi):

a linear function when you see those words think "line" as in y= mx+b notice that y depends on x. (to get y's value, you sub in a value for x and do the arithmetic) so part (a) Celsius as a linear function of degrees Fahrenheit means treat Celsius as y ( it depends on Fahrenheit) and Fahrenheit as x You have two (x,y) points: (32,0) and (212,100) can you find the equation of a line through two points? at the end, relabel x as F and y as C

OpenStudy (anonymous):

so I graph (32,0) and (212,100)?

OpenStudy (phi):

the question is asking for an equation. Normally you would use the 2 point formula, or you could first find the slope (change in y divided by change in x) and then use y-y1 = m(x-x1) where (x1,y1) is one of the points they give you.

OpenStudy (anonymous):

so it would be 100-32= m(212-0)

OpenStudy (phi):

almost. First find the slope.

OpenStudy (anonymous):

how do u do that?

OpenStudy (phi):

when you have time watch http://www.khanacademy.org/math/algebra/linear-equations-and-inequalitie/v/point-slope-and-standard-form and http://www.khanacademy.org/math/algebra/linear-equations-and-inequalitie/v/slope-of-a-line but slope is the difference between the 2 y values (subtract them) divided by the difference in the x values)

OpenStudy (anonymous):

oh ok i will thank u i think i got it from here :)

OpenStudy (anonymous):

@phi the slope is 5/9

OpenStudy (phi):

yes.

OpenStudy (phi):

for (a) you want y= 5/9 x + b you need to find b one way is replace x and y from one of the points (pick the one with zero, it's easier)

OpenStudy (anonymous):

ok i will do that right now

OpenStudy (anonymous):

so I replace x and y with zero and do the problem?

OpenStudy (phi):

replace (x,y) with (32,0) x=32 y=0

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

0=160/9+b then do i -160/9 from both sides?

OpenStudy (phi):

yes

OpenStudy (anonymous):

160/9=b 17.7=b

OpenStudy (anonymous):

-160/9 -17.7*

OpenStudy (phi):

ok, except I would leave it as -160/9 (it makes the equation exact)

OpenStudy (anonymous):

oh ok thank u :) what do i do for c? i dont understand what im supposed to do

OpenStudy (phi):

first, what did you get for (a)?

OpenStudy (anonymous):

a. i got -160/9 right? or was there more to it?

OpenStudy (phi):

-160/9 is b you want to whole equation y= mx+b for (a)

OpenStudy (anonymous):

how do you solve for that ... like this? 100=5/9(212)+(-160/90

OpenStudy (anonymous):

(-160/9)*

OpenStudy (phi):

that will be true. but all (x,y) pairs will be true. make y and x variables in the equation

OpenStudy (phi):

the whole point is we want an equation where you put in an x, and the equation tells you how to find y.

OpenStudy (anonymous):

y=5/9(212)+(-160/9)

OpenStudy (phi):

close. but the 212 should be x you get y= (5/9)x -160/9 now if you put x= 212, you can find y: 5*212/9 -160/9= 100 . We find 212 F = 100 C!

OpenStudy (phi):

So for (a) I would replace x and y with F and C, so it's easier to understand: C= (5/9)F - 160/9

OpenStudy (anonymous):

100=(5/9)(212)-160/9 so the y-intercept is c? which is 100

OpenStudy (phi):

Better go watch the videos (search for y intercept). But in the equation y= mx+b m is the slope and b is the y-intercept. for this equation y= (5/9)x - 160/9 m= 5/9 and b (the y-intercept) is -160/9

OpenStudy (phi):

so (b) the slop = 5/9. It means that 1 degree change in Fahrenheit is only 5/9 of a degree change in Celsius. (Fahrenheit degrees are smaller)

OpenStudy (anonymous):

oh ok b=y-intercept

OpenStudy (phi):

for (c) the y-intercept is -160/9 or about -17.8 this means that when it is 0 F it is -17.8 C or about -18 C when 0 F

OpenStudy (phi):

oh ok b=y-intercept that is why they call y= mx+b "slope intercept form" m =slope, b =y-intercept

OpenStudy (anonymous):

oh i see now thank you for all of your help! :) ill go watch the videos now!

OpenStudy (phi):

They are pretty good if you have enough popcorn.

OpenStudy (anonymous):

lol awesome